OpenStudy (anonymous):

The formula for degrees Celsius is: C = 9 over 5 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 4 degrees Celsius is.

OpenStudy (anonymous):

formula for f multiply by 5(F-32) divide by 5C add 32

OpenStudy (anonymous):

plug 4 into the original equation

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

That's wrong :( hold on... :(

OpenStudy (anonymous):

C = (5/9) (F - 32) So solve for F by first multiplying both side by (9/5): (9/5)C = (9/5)(5/9)(F-32) (9/5)C = F - 32 Then add 32 to both sides: (9/5)C + 32 = F So: F = (9/5)C + 32